P is the resultant of A and B. Q is resultant of A and (-B), the value of P² + Q² would be
(1) 2B²
(2) 2A²
(3) 2(A²-B²)
(4) 2(A²+B²)

Grade:11

1 Answers

Mahima Kanawat
5 years ago
Dear student 
P=A+B
Q=A-B
P2+Q= (A+B) 2+ (A-B) = A2+B2+2AB+A2+B2-2AB
=2(A2+B2)
WITH REGARDS 
MAHIMA 
ASKIITIANS FORUM EXPERT